OpenStudy (anonymous):

What is the DG under standard temperature for the following reaction between? NH3(s) -> NH3(l) Given: NH3(s): DHf = -46.11 kJ; DSf = 192.45 J/K NH3(l): DHf = -80.29 kJ; DSf = 111.3 J/K

OpenStudy (jfraser):

find the DH of the reaction, find the DS of the reaction, then plug into DG=DH-TDS

OpenStudy (anonymous):

would the DH be -34.18?

OpenStudy (anonymous):

I am sorry but I can't help with this one sorry/ :(

OpenStudy (jfraser):

DH, DS, and DG are always products - reactants, so is that the right order?

OpenStudy (anonymous):

@JFraser I think it is (-80.29 kJ)-(-46.11)= -34.18

OpenStudy (jfraser):

so that's the DH, do the same thing to find the DS

OpenStudy (anonymous):

I have that too, DS=-81.15

OpenStudy (anonymous):

What I don't know is how to set up or plug in the info into the equation DG=DH-TDS

OpenStudy (jfraser):

you just found DH and DS, what's the temperature?

OpenStudy (anonymous):

well.. It says to use the standard temperature, so that would be 273.15 K..?

OpenStudy (jfraser):

correct, so plug all 3 pieces into \[\Delta G = \Delta H - T * \Delta S\] and be sure to convert the KJ and J into the same unit

OpenStudy (anonymous):

okay, Thank you @JFraser